# Synthesizer: a Software Package for Synthetic Astronomical Observables

### v2

@@ Abstract
We present Synthesizer, a fast, flexible, modular and extensible platform for modelling synthetic astrophysical observables. Synthesizer can be used for a number of applications, but is predominantly designed for generating mock observables from analytical and numerical galaxy formation simulations. These use cases include (but are not limited to) analytical modelling of the star formation and metal enrichment histories of galaxies, the creation of mock images and integral field unit observations from particle based simulations, detailed photoionisation modelling of the central regions of active galactic nuclei, and spectro-photometric fitting. We provide a number of stellar population synthesis models, photoionisation code configurations, dust models, and imaging configurations that can be used ‘out-of-the-box’ interactively. We invite and encourage the community to use, test and develop the code, and hope that the foundation developed will provide a flexible framework for a number of tasks in forward and inverse modelling of astrophysical observables. The code is publicly available at synthesizer-project.github.io.
@@ 1 Introduction
There are many ways of comparing theoretical models of galaxy formation and evolution with observations. One of the most common approaches is to fit the observations to derive constraints on resolved and integrated physical properties and parameters (Walcher et al., 2011; Conroy, 2013). This procedure, known as spectral energy distribution (SED) fitting, or ‘inverse modelling’, involves many simplifying assumptions (Pacifici et al., 2023), which can lead to biases in derived parameters, such as the galaxy stellar mass (e.g. Lee et al., 2009; Mitchell et al., 2013; Hayward & Smith, 2015; Lower et al., 2020). Despite these uncertainties, the derived properties are often used to test phenomenological models, or compared directly with the predictions of numerical simulations in this physical properties space (Somerville & Davé, 2015).
@@
The emission from co-eval stellar populations is typically modelled by assuming an initial mass function (IMF) and a stellar population synthesis (SPS) model, that can include a range of different modelling assumptions, such as stellar multiplicity (Eldridge et al., 2017; Stanway & Eldridge, 2018a; Byrne et al., 2022), or the contribution of rare stellar phases (e.g. Maraston, 2005). The aggregate SFZH of the galaxy is then a composite of these co-eval populations. This can be modelled using analytic prescriptions, which provide the integrated SFZH in a parametric (or functional) form (Iyer & Gawiser, 2017; Carnall et al., 2019), or through more flexible non-parametric (or non-functional) approaches (Leja et al., 2019). Semi-analytic models, based on dark matter only numerical simulations, can model more complex star formation histories (Baugh, 2006), including the effects of galaxy mergers, but do not include spatial information (except in a simplified form in more recent models, Henriques et al., 2020; Stevens et al., 2024). These simulations have been used to forward model the galaxy population in large periodic volumes (Kitzbichler & White, 2007; Henriques et al., 2011; Somerville et al., 2012). Cosmological hydrodynamic simulations explicitly model the 3D star-gas-black hole geometry, with detailed merger and assembly histories (Crain & van de Voort, 2023), and each stellar element (or particle) can be treated as an individual co-eval population.
@@
Finally, a number of codes exist for transforming the forward modelled outputs into data products that closely resemble those produced from observational instruments (Harborne et al., 2023; Gabrielpillai et al., 2024). A number of codes wrap these various aspects into full pipelines, that can generate field-level observables (Fortuni et al., 2023; Marshall et al., 2024).
@@
In addition, new accelerated methods utilising neural density estimators, known as Simulation Based Inference (SBI) methods, have emerged in recent years (Cranmer et al., 2020; Ho et al., 2024). These methods require large synthetic training sets, and can easily accommodate wide sets of parameter variations, to test degeneracies in inference schemes. Generating such large training sets requires rapid forward models of the integrated and resolved galaxy emission, as well as field-level instrumental effects (Lovell et al., 2024; Fischbacher et al., 2024a, b; Çakır et al., 2024). Population level models, such as pop-cosmos (Alsing et al., 2024), also require rapid forward models in order to train emulators (Alsing et al., 2020), allowing hierarchical inference throughout the forward model.
@@ 2 Design Overview
Figure 2: An example showing some of the forward modelled observables that can be produced using Synthesizer. We show a galaxy taken from the Illustris TNG50 simulation, processed using a line of sight dust attenuation model, assuming a fixed dust to metal ratio. Top row: full colour RGB images in, from left to right, JWST/NIRCam [F115W,F150W,F200W], HST/WFC3 [F450W,F547M,F814W], and Euclid NISP [YJH], where each band is linearly combined. The NIRCam image is produced using a PSF taken from Perrin et al. (2014). Middle row, left to right: stellar mass map, mass-normalised metallicity map, mass-normalised age map, optical depth map (obtained from the gas particles), H-$\alpha$ emission map. Bottom row, left to right: the star formation - metal enrichment history (with marginal age and metallicity distributions), intrinsic and dust attenuated spectra and derived photometry (for edge on and face on orientations), histogram of optical depth for face-on and edge-on line of sight calculations, BPT diagram (star shows the integrated ratios, points show individual star particles, lines show the Kauffmann & Haehnelt (2000); Kewley et al. (2013) classification regions).
@@ 3.2 Gas Components
A Gas component requires at minimum the particle masses and metallicities. Gas components can also additionally contain information such as the coordinates, smoothing lengths, dust masses and / or the dust to metal (DtM) ratios, necessary for calculating the LoS attenuation. If the DtM ratio is not provided then a default $\rm DtM=0.3$ is assumed. If dust masses are not provided then this DtM is used to automatically calculate the dust masses. Other ‘subgrid’ properties of gas particles can also be defined, such as the HI and HII mass fractions.
@@ 4 Grids
Figure 5: Example of the spectra contained within a grid object. Top left: all the different spectra produced from a grid, with age 5 Myr and metallicity 0.01, using the BC03 SPS model with a Chabrier IMF, and post-processed with Cloudy with the fiducial photoionisation parameters. Top right: multiple incident grids assuming different SPS models (BC03, BPASS, FSPS, M24) all at the same age (5 Myr) and metallicity ($Z=0.01$) Bottom left: variation of the same model in the top left panel with age, at fixed metallicity ($Z=0.01$). Bottom right: variation of the same model in the top left panel with metallicity, at fixed age (5 Myr).
@@ 4.1 Stellar Population Synthesis Grids
specifies that the grid is constructed using v2.2.1 of the Binary Population and Spectral Synthesis (BPASS) SPS model for the binary (bin) variant, assumes a Chabrier (2003) IMF between 0.1 and 300 Msol, and that photoionisation modelling is performed using v17.03 of the Cloudy photoionisation code assuming our default assumptions (more details below). Many grids will omit parameters in the name where they take the default values. It is also worth noting that all grids files contain attributes describing the parameters used in detail, with units, and these should be the values used when referencing programmatically.
@@
Further details on each model are provided in the listed references. For the isochrones, Padova refers to the collection of isochrones by Alongi et al. (1993); Bressan et al. (1993); Fagotto et al. (1994a, b); Girardi et al. (1996, 2000) and the Geneva isochrones refer to the compilation by Schaller et al. (1992); Charbonnel et al. (1996, 1999). For information on the MIST and BPASS isocontours, see Dotter (2016); Choi et al. (2016) and Section 2 of Stanway & Eldridge (2018b), respectively.
@@ 4.3.1 SPS Grids
For a fixed values of $\langle U\rangle$ and $Q$, the geometry of the region is encapsulated in the $\epsilon^{2}n_{\mathrm{H}}$ term. The advantage of using $\langle U\rangle$ is that it significantly reduces the dimensionality of our input grid: any combination of $Q$, $\epsilon^{2}$ and $n_{\mathrm{H}}$ that yields the same $\langle U\rangle$ will result in an identical nebular spectrum – provided the shape of the ionising spectrum and metallicity are held constant.
@@ 6 Generating Spectra
Figure 8: Composite galaxy spectra example. The stellar component has a mass of $10^{11}\,{\rm M_{\odot}}$, and assumes a log-normal star formation history ($\tau=1$, peak age = 400 Myr, max age = 2 Gyr) and solar metallicity ($Z=0.01$). The emission is obtained using the Pacman model, assuming an escape fraction, $f_{\rm esc}=0.1$, using the BPASS v2.3 SPS model with the fiducial photoionisation parameters. The AGN component assumes a NLR / BLR covering fraction of 0.1, ionisation parameter of 0.1, and a Hydrogen density of $10^{5}$ cm-3. Both components are attenuated by a diffuse ISM component assuming a MWN18 attenuation curve, with $\tau_{\rm V}=0.33$. The young ($\leqslant 10\,{\rm Myr}$) stellar component is further attenuated by a factor of 0.67. Dust emission is modelled using a Greybody, with temperature 30 K and emissivity of 1.2.
@@ 6.1 Particle Based Models
In both schemes, the spectra assigned to a given particle depends on whether the grid dimensions are described linearly or logarithmically. Users can specify this choice for each dimension as required, however recommended defaults are saved in each grid file. For age and metallicity we recommend using logarithmic scaling.
@@ 7.4 Photometry
where $f_{\lambda}$ is the wavelength dependent flux density incident on the filter, in units of $\rm erg/s/cm^{2}/\textup{\leavevmode\nobreak\ \AA }$}. For frequency dependent spectral flux density we convert $f_{\nu}=\lambda f_{\lambda}/c^{2}$. The mean of the flux density is then given by
$\overline{f_{\lambda}}(T)\,/\,({\rm erg\,s^{-1}\,cm^{-2}\,\textup{\leavevmode\nobreak\ \AA }^{-1}})=\frac{\int_{\lambda}\lambda f_{\lambda}T(\lambda)d\lambda}{\int_{\lambda}\lambda T(\lambda)d\lambda}$ (8)
@@ 10 Summary & Conclusions
Auto differentiability is rapidly being implemented in a number of astrophysical codes to allow for gradient based inference and optimisation schemes. Whilst the class based structure of Synthesizer does not lend itself naturally to such approaches, we hope to explore the implementation of autodiff in various parts of the code to allow for these schemes.
@@ Appendix C Creating Your Own Grids
Advanced users can create their own Synthesizer grids. These can be intrinsic grids of stellar emission, generated from stellar population synthesis models, or grids post-processed through photoionisation codes such as cloudy. The code for creating custom grids is contained in an independent repository. [^15] https://github.com/synthesizer-project/grid-generation You will need a working installation of Synthesizer for these scripts to work, as well as other dependencies for specific codes (e.g. Cloudy, python-FSPS). Grids should follow the naming convention where possible, see Section 4.1. Please see Appendix B for details on how to modify the chemical abundance pattern of gas, stars and dust using the ‘abundances’ object, and use this when running cloudy.
@@ Appendix F Fiducial Emission Models
F.0.1 Screen Attenuation Model
@@
F.0.2 Charlot & Fall 2000 Attenuation Model
@@ Appendix H Parametric Stellar morphologies
Different parametric morphology components can be added together, to mimic the contribution of different stellar populations that are distinct spatially; for example, a bulge and disk can be combined to mimic a classical spiral galaxy. This can be achieved using the overridden addition operator on two image collections. An example of this is shown in Figure 13, showing a disk and bulge morphology, after coupling to a SFZH and grid, and how these can be used to generate RGB images from multi-band photometry.
@@ Appendix I Property Maps
The Line of Sight (LoS) attenuation model is a more sophisticated model for the diffuse interstellar dust attenuation that can be computed for a Particle Galaxy object, by taking into account the full three dimensional star-dust or black hole-dust geometry. Where the mass and distribution of dust is self-consistently predicted by the simulation (e.g. Aoyama et al., 2017; McKinnon et al., 2017; Gjergo et al., 2018; Li et al., 2019) these values can be used directly to calculate the dust column density towards each star or black hole particle, assuming a given viewing direction. Where this information is not available, the distribution of gas-phase metals can be used as a proxy for the dust distribution (e.g. Wilkins et al., 2017; Vijayan et al., 2021). In this approach the column density of gas-phase metals along the line-of-sight to each emitting particle is computed by summing the contributions from each gas particle, assuming a 3D kernel function. This column density is then related to the dust column by assuming some Dust-to-Metal (DtM) ratio. In both cases this approach requires details on the star particle distribution as well as the gas / dust. Subgrid dust e.g. within star forming clouds, can be accounted for within each star particle source.

### v3

@@ Abstract
We present Synthesizer, a fast, flexible, modular and extensible platform for modelling synthetic astrophysical observables. Synthesizer can be used for a number of applications, but is predominantly designed for generating mock observables from analytical and numerical galaxy formation simulations. These use cases include (but are not limited to) analytical modelling of the star formation and metal enrichment histories of galaxies, the creation of mock images and integral field unit observations from particle based simulations, detailed photoionisation modelling of the central regions of active galactic nuclei, and spectro-photometric fitting. We provide a number of stellar population synthesis models, photoionisation code configurations, dust models, and imaging configurations that can be used ‘out-of-the-box’ interactively. The code can be used to quantitatively test the dependence of forward modelled observables on various model and parameter choices, and rapidly explore large parameter ranges for calibration and inference tasks. We invite and encourage the community to use, test and develop the code, and hope that the foundation developed will provide a flexible framework for a number of tasks in forward and inverse modelling of astrophysical observables. The code is publicly available at synthesizer-project.github.io.
@@ 1 Introduction
There are many ways of comparing theoretical models of galaxy formation and evolution with observations. One of the most common approaches is to fit the observations to derive constraints on resolved and integrated physical properties and parameters (Walcher et al., 2011; Conroy, 2013). This procedure, known as spectral energy distribution (SED) fitting, or ‘inverse modelling’, involves many simplifying assumptions (Pacifici et al., 2023), such as parameterised star formation histories, and screen models for dust attenuation, which can lead to biases in derived parameters, such as the galaxy stellar mass (e.g. Lee et al., 2009; Mitchell et al., 2013; Hayward & Smith, 2015; Lower et al., 2020). Despite these uncertainties, the derived properties are often used to test phenomenological models, or compared directly with the predictions of numerical simulations in this physical properties space (Somerville & Davé, 2015).
@@
The emission from co-eval stellar populations is typically modelled by assuming an initial mass function (IMF) and a stellar population synthesis (SPS) model, that can include a range of different modelling assumptions, such as stellar multiplicity (Eldridge et al., 2017; Stanway & Eldridge, 2018a; Byrne et al., 2022), or the contribution of rare stellar phases (e.g. Maraston, 2005). The aggregate star formation and metal enrichment history (SFZH) of the galaxy is then a composite of these co-eval populations. This can be modelled using analytic prescriptions, which provide the integrated SFZH in a parametric (or functional) form (Iyer & Gawiser, 2017; Carnall et al., 2019), or through more flexible non-parametric (or non-functional) approaches (Leja et al., 2019). Semi-analytic models, based on dark matter only numerical simulations, can model more complex star formation histories (Baugh, 2006), including the effects of galaxy mergers, but do not include spatial information (except in a simplified form in more recent models, Henriques et al., 2020; Stevens et al., 2024). These simulations have been used to forward model the galaxy population in large periodic volumes (Kitzbichler & White, 2007; Henriques et al., 2011; Somerville et al., 2012). Cosmological hydrodynamic simulations explicitly model the 3D star-gas-black hole geometry, with detailed merger and assembly histories (Crain & van de Voort, 2023), and each stellar element (or particle) can be treated as an individual co-eval population.
@@
Finally, a number of codes exist for transforming the forward modelled outputs into data products that closely resemble those produced from observational instruments (Harborne et al., 2023; Gabrielpillai et al., 2024). A number of codes wrap these various aspects into full pipelines, that can generate field-level observables (Fortuni et al., 2023; Marshall et al., 2025).
@@
In addition, new accelerated methods utilising neural density estimators, known as Simulation Based Inference (SBI) methods, have emerged in recent years (Cranmer et al., 2020; Ho et al., 2024). These methods require large synthetic training sets, and can easily accommodate wide sets of parameter variations, to test degeneracies in inference schemes. Generating such large training sets requires rapid forward models of the integrated and resolved galaxy emission, as well as field-level instrumental effects (Lovell et al., 2024; Fischbacher et al., 2024, 2025; Çakır et al., 2024). Population level models, such as pop-cosmos (Alsing et al., 2024), also require rapid forward models in order to train emulators (Alsing et al., 2020), allowing hierarchical inference throughout the forward model.
@@ 2 Design Overview
Figure 2: An example showing some of the forward modelled observables that can be produced using Synthesizer. We show a galaxy taken from the Illustris TNG50 simulation, processed using a line of sight dust attenuation model, assuming a fixed dust to metal ratio, and using the BPASS v2.2.1 SPS model, with binary stars and a Kroupa IMF. Top row: full colour RGB images in, from left to right, JWST/NIRCam [F115W,F150W,F200W], HST/WFC3 [F450W,F547M,F814W], and Euclid NISP [YJH], where each band is linearly combined. The NIRCam image is produced using a PSF taken from Perrin et al. (2014). Middle row, left to right: stellar mass map, mass-normalised metallicity map, mass-normalised age map, optical depth map (obtained from the gas particles), H-$\alpha$ emission map. Bottom row, left to right: the star formation - metal enrichment history (with marginal age and metallicity distributions), intrinsic and dust attenuated spectra and derived photometry (for edge on and face on orientations), histogram of optical depth for face-on and edge-on line of sight calculations, BPT diagram (star shows the integrated ratios, points show individual star particles, lines show the Kauffmann & Haehnelt (2000); Kewley et al. (2013) classification regions).
@@ 3.2 Gas Components
A Gas component requires at minimum the particle masses and metallicities. Gas components can also additionally contain information such as the coordinates, smoothing lengths, dust masses and / or the dust to metal (DtM) ratios, necessary for calculating the LoS attenuation. If the DtM ratio is not provided then a default $\rm DtM=0.3$ is assumed (Camps et al., 2016; Schulz et al., 2020). If dust masses are not provided then this DtM is used to automatically calculate the dust masses. Other ‘subgrid’ properties of gas particles can also be defined, such as the HI and HII mass fractions.
@@ 4 Grids
Figure 5: Example of the spectra contained within a grid object. Top left: all the different spectra produced from a grid, with age 5 Myr and metallicity 0.01, using the BPASS v2.2.1 SPS binary model with a Kroupa IMF, and post-processed with Cloudy with the fiducial photoionisation parameters. Top right: multiple incident grids assuming different SPS models (BC03, BPASS, FSPS, M24) all at the same age (5 Myr) and metallicity ($Z=0.01$) Bottom left: variation of the same model in the top left panel with age, at fixed metallicity ($Z=0.01$). Bottom right: variation of the same model in the top left panel with metallicity, at fixed age (5 Myr).
@@ 4.1 Stellar Population Synthesis Grids
specifies that the grid is constructed using v2.2.1 of the Binary Population and Spectral Synthesis (BPASS) SPS model for the binary (bin) variant, assumes a Chabrier (2003) IMF between 0.1 and 300 $M_{\odot}$, and that photoionisation modelling is performed using v17.03 of the Cloudy photoionisation code assuming our default assumptions (more details below). Many grids will omit parameters in the name where they take the default values. It is also worth noting that all grids files contain attributes describing the parameters used in detail, with units, and these should be the values used when referencing programmatically.
@@
Further details on each model are provided in the listed references. For the isochrones, Padova refers to the collection of isochrones by Alongi et al. (1993); Bressan et al. (1993); Fagotto et al. (1994a, b); Girardi et al. (1996, 2000) and the Geneva isochrones refer to the compilation by Schaller et al. (1992); Charbonnel et al. (1996, 1999). For information on the MIST and BPASS isocontours, see Dotter (2016); Choi et al. (2016) and Section 2 of Stanway & Eldridge (2018b), respectively. Yggdrasil is an SPS model specifically designed for Population III stars, and as such does not have a typical grid metallicity structure due to the predominance of metal free models.
@@ 4.3.1 SPS Grids
For fixed values of $\langle U\rangle$ and $Q$, the geometry of the region is encapsulated in the $\epsilon^{2}n_{\mathrm{H}}$ term. The advantage of using $\langle U\rangle$ is that it significantly reduces the dimensionality of our input grid: any combination of $Q$, $\epsilon^{2}$ and $n_{\mathrm{H}}$ that yields the same $\langle U\rangle$ will result in an identical nebular spectrum – provided the shape of the ionising spectrum and metallicity are held constant.
@@ 6 Generating Spectra
Figure 8: Composite galaxy spectra example. The stellar component has a mass of $10^{11}\,{\rm M_{\odot}}$, and assumes a log-normal star formation history ($\tau=1$, peak age = 400 Myr, max age = 2 Gyr) and solar metallicity ($Z=0.014$). The emission is obtained using the Pacman model, assuming an escape fraction, $f_{\rm esc}=0.1$, using the BPASS v2.3 SPS model with the fiducial photoionisation parameters. The AGN component assumes a NLR / BLR covering fraction of 0.1, ionisation parameter of 0.1, and a Hydrogen density of $10^{5}$ cm-3. Both components are attenuated by a diffuse ISM component assuming a MWN18 attenuation curve, with $\tau_{\rm V}=0.33$. The young ($\leqslant 10\,{\rm Myr}$) stellar component is further attenuated by a factor of 0.67. Dust emission is modelled using a Greybody, with temperature 30 K and emissivity of 1.2.
@@ 6.1 Particle Based Models
In both schemes, the spectra assigned to a given particle depends on whether the grid dimensions are described linearly or logarithmically. Users can specify this choice for each dimension as required, however recommended defaults are saved in each grid file. For age and metallicity we recommend using logarithmic scaling. For all grids, we truncate the predictions to the edge of the grid, and do not extrapolate beyond the grid boundaries; users should take care that either their input stellar properties are covered by the chosen grid, or that they trust the predictions at the grid boundaries for out of distribution particles. This is also the case for parametric models, described below.
@@ 7.4 Photometry
where $f_{\lambda}$ is the wavelength dependent flux density incident on the filter, in units of $\rm erg/s/cm^{2}/\textup{\penalty 10000\ \AA }$}. For frequency dependent spectral flux density we convert $f_{\nu}=\lambda f_{\lambda}/c^{2}$. The mean of the flux density is then given by
$\overline{f_{\lambda}}(T)\,/\,({\rm erg\,s^{-1}\,cm^{-2}\,\textup{\penalty 10000\ \AA }^{-1}})=\frac{\int_{\lambda}\lambda f_{\lambda}T(\lambda)d\lambda}{\int_{\lambda}\lambda T(\lambda)d\lambda}$ (8)
@@ 10 Summary & Conclusions
A code like Synthesizer has many evident use cases. Users that wish to forward model key observable distribution functions from cosmological simulations, such as luminosity functions and colours, can now do so for different SPS, dust and photoionisation models, simply by changing a few small parameters in the pipeline. This unlocks the possibility of robustly assessing the dependence of the predictions on the forward model, providing quantitative baseline uncertainties on these quantities, that can be compared to the uncertainties from inverse modelling approaches. The inclusion of AGN and stellar emission in a single unified code also opens up the possibility to assess the relative contribution of these components to galaxy emission at different wavelengths and across cosmic time. Finally, the speed of Synthesizer opens up a number of opportunities in Simulation Based Inference (SBI) schemes; an example of this has already been demonstrated in (Lovell et al., 2024), by applying to the large suite of CAMELS simulations. Numerous other applications of Synthesizer exist that we cannot address here due to space limitations; we look forward to seeing the community explore both these applications and creative uses the authors have not yet envisaged.
@@
Auto differentiability is rapidly being implemented in a number of astrophysical codes to allow for gradient based inference and optimisation schemes (e.g. Pandya et al. in prep.). Whilst the class based structure of Synthesizer does not lend itself naturally to such approaches, we hope to explore the implementation of autodiff in various parts of the code to allow for these schemes.
@@ Appendix C Creating Your Own Grids
Advanced users can create their own Synthesizer grids. These can be intrinsic grids of stellar emission, generated from stellar population synthesis models, or grids post-processed through photoionisation codes such as cloudy. The code for creating custom grids is contained in an independent repository, https://github.com/synthesizer-project/grid-generation. You will need a working installation of Synthesizer for these scripts to work, as well as other dependencies for specific codes (e.g. Cloudy, python-FSPS). Grids should follow the naming convention where possible, see Section 4.1. Please see Appendix B for details on how to modify the chemical abundance pattern of gas, stars and dust using the ‘abundances’ object, and use this when running cloudy.
@@ Appendix F Fiducial Emission Models
F.1 Screen Attenuation Model
@@
F.2 Charlot & Fall 2000 Attenuation Model
@@ Appendix H Parametric Stellar morphologies
Different parametric morphology components can be added together, to mimic the contribution of different stellar populations that are distinct spatially; for example, a bulge and disk can be combined to mimic a classical spiral galaxy. This can be achieved using the overridden addition operator on two image collections. An example of this is shown in Figure 13, showing a disk and bulge morphology, after coupling to a star formation and metal enrichment history (SFZH) and grid, and how these can be used to generate RGB images from multi-band photometry.
@@ Appendix I Property Maps
The Line of Sight (LoS) attenuation model is a more sophisticated model for the diffuse interstellar dust attenuation that can be computed for a Particle Galaxy object, by taking into account the full three dimensional star-dust or black hole-dust geometry. Where the mass and distribution of dust is self-consistently predicted by the simulation (e.g. Aoyama et al., 2017; McKinnon et al., 2017; Gjergo et al., 2018; Li et al., 2019) these values can be used directly to calculate the dust column density towards each star or black hole particle, assuming a given viewing direction. Where this information is not available, the distribution of gas-phase metals can be used as a proxy for the dust distribution (e.g. Wilkins et al., 2017; Vijayan et al., 2021). In this approach the column density of gas-phase metals along the line-of-sight to each emitting particle is computed by summing the contributions from each gas particle, assuming a 3D kernel function. This column density is then related to the dust column by assuming some Dust-to-Metal (DtM) ratio. This DtM can be fixed, or depend on integrated properties of the galaxy, such as the gas phase metallicity (Rémy-Ruyer et al., 2014) or redshift (Vogelsberger et al., 2020). In both cases this approach requires details on the star particle distribution as well as the gas / dust. Subgrid dust e.g. within star forming clouds, can be accounted for within each star particle source.